# Corona rigidity

### v3

@@ 5.4. $\mathrm{C}^{*}$-algebras
for all $x$ and $y$ in the domain, and proved that for such a function $f$ between abelian $\mathrm{C}^{*}$-algebras there is a -homomorphism $\Phi$ such that $\|\Phi-f\|\leq((4+\frac{3}{2}(1+\pi))\varepsilon+1-\sqrt{1-4\varepsilon})\|f\|$ ([140, Theorem 5.1]).
@@ 5.5. Reduced products of $\mathrm{C}^{*}$-algebras
Is the category of $\mathrm{C}^{*}$-algebras of the form $C(X,M_{n}(\mathbb{C}))$, where $X$ is compact and metrizable, Ulam-stable?
@@ 5.6. The Kadison–Kastler conjecture
Since our main interest is in isomorphisms, it is natural to consider the following notion ([110, Definition 1.2]). Recall that by $A_{1}$ we denote the unit ball of a $\mathrm{C}^{*}$-algebra $A$.
@@
We say that $A$ satisfies the Kadison–Kastler conjecture if such an $\varepsilon$ exists. A class $\mathcal{C}$ of $\mathrm{C}^{*}$-algebras is said to be Kadison–Kastler stable (or KK stable) if each algebra in $\mathcal{C}$ satisfies the Kadison–Kastler conjecture and the same $\varepsilon$ witnesses stability uniformly on elements of $\mathcal{C}$. The class of nonseparable $\mathrm{C}^{*}$-algebras is not KK stable, and there is even a nonseparable $\mathrm{C}^{*}$-algebra that does not satisfy the Kadison–Kastler conjecture ([27], see also [59, §14.4]). The class of unital separable amenable [^16] Amenable, or nuclear, $\mathrm{C}^{*}$-algebras form the most studied and arguably the most important class of $\mathrm{C}^{*}$-algebras. $\mathrm{C}^{*}$-algebras is KK stable ([28]), but it is not known whether the class of all separable unital $\mathrm{C}^{*}$-algebras is KK stable.
@@
In this section we analyze constructions of isomorphisms of Borel quotient structures which are not topologically trivial, therefore giving partial answers to Question 1.1. To do this, we often rely on additional set theoretic assumptions such as the Continuum Hypothesis or its weakenings.
@@ 6.1.1. $\sigma$-complete back-and-forth systems
Each isomorphism between metric structures of density character $\aleph_{1}$ corresponds to a $\sigma$-complete back-and-forth system (see [59, Proposition 16.6.1]). If $A$ and $B$ are metric structures, a $\sigma$-complete back-and-forth system between $A$ and $B$ is a poset ${\mathbb{F}}$ such that each element $p$ of ${\mathbb{F}}$ denotes a partial isomorphism $\Phi^{p}:A^{p}\to B^{p}$ for $A^{p}\subseteq A$ and $B^{p}\subseteq B$, with the following properties:
@@
A structure is countably saturated if every consistent type over a countable set is realized.
Thus ‘countably saturated’ is what is usually called ‘$\aleph_{1}$-saturated’ and should not be confused with the notion of ‘countably saturated’ as defined in e.g., [25, §2.3] (such models are countable by definition, and countably saturated models in our sense are necessarily uncountable). The reason for this change of terminology is the same as the reason why countable additivity of Lebesgue measure is rarely referred to as $\aleph_{1}$-additivity. If the language is countable, then countable saturation is equivalent to countable compactness, i.e., the requirement that every countable consistent type is realized.
If two structures are countably saturated and elementarily equivalent, then there is a $\sigma$-complete back-and-forth system between them (see [59, Theorem 16.6.4]). This implies a special case of the following theorem due to Keisler (e.g., [25, Theorem 5.1.13]).
@@
The countable saturation of metric structures of density character $\aleph_{1}$ also has fundamental effects on the automorphism groups, since it allows the construction of $2^{\aleph_{1}}$ distinct $\sigma$-complete back-and-forth systems, giving rise to $2^{\aleph_{1}}$ distinct automorphisms of the structure (see [59, Theorem 16.6.3]).
@@ 6.1.2. Borel quotients of $\mathcal{P}({\mathbb{N}})$
However, in [95] CH was used to prove that the quotients over ${\mathcal{Z}}_{0}$, ${\mathcal{Z}}_{\log}$, and even all so-called EU-ideals, are isomorphic. The intricate back-and-forth construction produced an isometric isomorphism. It took a couple of decades before this heroic achievement was put into a proper context: All quotients of $\mathcal{P}({\mathbb{N}})$ by generalized density ideals (Definition 3.5) are countably saturated metric structures by Proposition 5.11 and Theorem 6.8. This fact was used in [54, Theorem 2] to prove that under CH there are only two non-isomorphic quotients over dense density ideals, and it implies the following.
@@
Assume $\mathrm{CH}$ and let $\mathcal{I}\subseteq\mathcal{P}({\mathbb{N}})$ be a generalized density ideal. Then there are $2^{\mathfrak{c}}$ topologically nontrivial automorphisms of $\mathcal{P}({\mathbb{N}})/\mathcal{I}$.
@@ 6.1.3. Zero-dimensional remainders
In [121] Parovičenko characterized, under $\mathrm{CH}$, $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ as the only space sharing certain topological properties (such properties define a Parovičenko space, in today’s terminology). These topological properties are indeed shared by all spaces of the form $\beta X\setminus X$, where $X$ is a second countable locally compact noncompact zero-dimensional space [^18] It was proved in [165] that $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ being unique among Parovičenko spaces is equivalent to $\mathrm{CH}$.. Through Stone duality, Parovičenko’s theorem essentially states that the Boolean algebras of clopen sets of those remainders are atomless and countably saturated. Since the theory of atomless Boolean algebras is complete, all of its models are elementarily equivalent. With the definition of algebraically trivial homeomorphisms (Definition 4.8) in mind and the fact that topologically trivial automorphisms of coronas of abelian $\mathrm{C}^{*}$-algebras are dual to algebraically trivial homeomorphisms ([170, Proposition 2.7]), this implies the following.
@@
On the positive side, if $X$ is equal to a strictly increasing union of compact subsets $K_{n}$ which satisfy ($\partial K_{n}$ is the topological boundary of $K_{n}$) $\sup_{n}|\partial K_{n}|<\infty$, then $C(\beta X\setminus X)$ is countably saturated (this is a special case of [71, Theorem 2.5], and includes the case $X=\mathbb{R}$).
@@ 6.1.4. Countable saturation of the reduced products
Our next task is to review some of the applications of Theorem 6.3 in the noncommutative setting of reduced products. Ultraproducts associated with countably incomplete ultrafilters are countably saturated (see [25, Theorem 6.1.1]). This, in particular, is the case for the norm ultraproducts of $\mathrm{C}^{*}$-algebras. As for coronas of separable $\mathrm{C}^{*}$-algebras, they are not necessarily countably saturated. Notably, the Calkin algebra is not countably saturated (§6.3). However, corona algebras that are reduced products over $\operatorname{Fin}$ are countably saturated by the following result ([71], see also [59, Theorem 16.5.1]).
@@ 6.1.5. Metric Feferman–Vaught theorem
Theorem 6.8 is proved by combining the metric version of the classical Feferman–Vaught theorem with the proof of the countable saturation of ultraproducts associated with countably incomplete ultrafilters (see e.g., [59, §16.5]). In classical model theory Feferman and Vaught ([75]) gave an effective way to determine the satisfaction of formulas in the reduced products of models of the same language associated with any ideal over ${\mathbb{N}}$. Their result was generalized in [80] to metric structures. The metric Feferman–Vaught theorem essentially states that if $(A_{n})_{n\in{\mathbb{N}}}$ are $\mathcal{L}$-structures and $\mathcal{I}$ is an ideal on ${\mathbb{N}}$ then the theory of the reduced product $\prod_{n}A_{n}/\bigoplus_{\mathcal{I}}A_{n}$ can be recursively computed from the theories of $A_{n}$ and the theory of $\mathcal{P}({\mathbb{N}})/\mathcal{I}$ as a Boolean algebra.
@@
An elementary embedding from $C(K,A)$ into $A_{\infty}$ was exhibited in [61, Proposition 3.5], and the general case follows from Theorem 6.10. ∎
@@ 6.1.6. The Palyutin–Palmgren approach
The key to Palmgren’s proof is the observation that in every reduced power associated with $\operatorname{Fin}$ (or any other ideal such that the quotient Boolean algebra is atomless) every formula is equivalent to a Boolean combination of $h$-formulas. Therefore every type over such reduced product is axiomatized by a set consisting of $h$-formulas and negations of $h$-formulas. This proof was extended to prove countable saturation of reduced powers over a larger class of the so-called layered ideals in [31, Theorem 4.1].
@@ 6.1.7. Asymptotic sequence algebras
For discrete structures ‘density character’ is of course a synonym for cardinality. For more see [61, Theorem A–F].
@@ 6.2. Full saturation: The dividing line
A result analogous to (6.2.1) for ultrapowers appears in [65, Theorem 5.6], and the original proof of (6.2.1) closely followed the proof of the latter. This proof was considerably simplified by using ideas of Palyutin and Palmgren (see §6.1.6). Item (6.2.2) is proven by injecting $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ into the quotient in a gap-preserving manner. Its proof applies in a wider context, showing for example that if $M$ is a reduced product over the asymptotic density zero ideal ${\mathcal{Z}}_{0}$ (see §3.1), or any other ${\mathcal{I}}$ such that $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is not countably saturated, whose theory is not stable then $M$ is not even countably saturated ([31, Corollary 2.3]).
@@ 6.3.1. The Fredholm index obstruction
The simplest example of the failure of countable saturation in the Calkin algebra is related to the original Brown–Douglas–Fillmore question, whether there is an automorphism of $\mathcal{Q}(H)$ that sends $\dot{s}$ (where $s$ is the unilateral shift) to its adjoint (see Question 4.16). This is really a question about orbits of unitaries in the Calkin algebra under the action of the automorphism group by conjugation. Every unitary $u$ in $\mathcal{Q}(H)$ lifts to a partial isometry (see (4.1.15)) between closed subspaces of $H$ of finite co-dimension. The difference between the co-dimensions of the domain and the range of the lift is its Fredholm index (§4.3). The function that sends $u$ to the Fredholm index of its lift is a group homomorphism from $\mathcal{U}(\mathcal{Q}(H))$ into $({\mathbb{Z}},+)$. This implies that the type of a unitary that (for example) has $2^{n}$th root for all $n$, but no $3$rd root, is consistent but not realized in $\mathcal{Q}(H)$. (The proof uses definability of the set of unitaries, in the sense of [9], see also [64, §3] to show that in $\mathcal{Q}(H)$—and in any $\mathrm{C}^{*}$-algebra—applying quantification over the unitary group to a definable predicate results in a definable predicate). By adding a variable for each $2^{n}$th root, this example provides a universal type which is consistent but not realized in $\mathcal{Q}(H)$.
@@ 6.3.2. Homotopy
By iterating the construction of Example 6.16 one obtains the following, due to N.C. Phillips ([63, Proposition 4.2]); recall the convention that $\dot{p}$ denotes a lift of $p$.
@@
There is a unital copy $A$ of the CAR algebra $\bigotimes_{\mathbb{N}}M_{2}(\mathbb{C})$ in $\mathcal{Q}(H)$ such that for every projection $p\in\mathcal{B}(H)$ with $\dot{p}\in A$, none of the tensor factors $M_{2}(\mathbb{C})$ lifts to a unital copy of $M_{2}(\mathbb{C})$ in $p\mathcal{B}(H)p$ (for more see [139]). Fix a unital copy $B$ of the CAR algebra that lifts to a unital copy of the CAR algebra in $\mathcal{B}(H)$ and an isomorphism $\Phi\colon A\to B$. The type of a unitary that implements this isomorphism is, by Example 6.16, consistent, but it is not realized in $\mathcal{Q}(H)$.
The Galois type of a tuple (i.e., its orbit under the action of the automorphism group) provides more information than its first-order type. Example 6.16 shows that in $\mathrm{C}^{*}$-algebras, the Galois type of a tuple needs to be supplemented with its homotopy type. There is an alternative way of defining the homotopy type of a tuple in a $\mathrm{C}^{*}$-algebra. For a definable set $X$ (such as the set of unitaries or the set of projections; see [64, §3]) one says that two elements of $X$ are homotopic if they can be connected by a continuous path inside $X$. This approach is used in defining K-theory groups of $\mathrm{C}^{*}$-algebras (see [131]), and we will take advantage of it in the next subsection.

### v4

@@ 5.4. $\mathrm{C}^{*}$-algebras
for all $x$ and $y$ in the domain, and proved that for such a function $f$ between abelian $\mathrm{C}^{*}$-algebras there is a -homomorphism $\Phi$ such that $\|\Phi-f\|\leq((4+\frac{3}{2}(1+\pi))\varepsilon+1-\sqrt{1-4\varepsilon})\|f\|$ ([146, Theorem 5.1]).
@@ 5.5. Reduced products of $\mathrm{C}^{*}$-algebras
Is the category of $\mathrm{C}^{*}$-algebras of the form $C(X,M_{n}(\mathbb{C}))$, where $X$ is compact and metrisable, Ulam-stable?
@@ 5.6. The Kadison–Kastler conjecture
Since our main interest is in isomorphisms, it is natural to consider the following notion ([116, Definition 1.2]). Recall that by $A_{1}$ we denote the unit ball of a $\mathrm{C}^{*}$-algebra $A$.
@@
We say that $A$ satisfies the Kadison–Kastler conjecture if such an $\varepsilon$ exists. A class $\mathcal{C}$ of $\mathrm{C}^{*}$-algebras is said to be Kadison–Kastler stable (or KK stable) if each algebra in $\mathcal{C}$ satisfies the Kadison–Kastler conjecture and the same $\varepsilon$ witnesses stability uniformly on elements of $\mathcal{C}$. The class of nonseparable $\mathrm{C}^{*}$-algebras is not KK stable, and there is even a nonseparable $\mathrm{C}^{*}$-algebra that does not satisfy the Kadison–Kastler conjecture ([27], see also [61, §14.4]). The class of unital separable amenable [^16] Amenable, or nuclear, $\mathrm{C}^{*}$-algebras form the most studied and arguably the most important class of $\mathrm{C}^{*}$-algebras. $\mathrm{C}^{*}$-algebras is KK stable ([28]), but it is not known whether the class of all separable unital $\mathrm{C}^{*}$-algebras is KK stable.
@@
In this section we analyse constructions of isomorphisms of Borel quotient structures which are not topologically trivial, therefore giving partial answers to Question 1.1. To do this, we often rely on additional set theoretic assumptions such as the Continuum Hypothesis or its weakenings.
@@ 6.1.1. $\sigma$-complete back-and-forth systems
Each isomorphism between metric structures of density character $\aleph_{1}$ corresponds to a $\sigma$-complete back-and-forth system (see [61, Proposition 16.6.1]). If $A$ and $B$ are metric structures, a $\sigma$-complete back-and-forth system between $A$ and $B$ is a poset ${\mathbb{F}}$ such that each element $p$ of ${\mathbb{F}}$ denotes a partial isomorphism $\Phi^{p}:A^{p}\to B^{p}$ for $A^{p}\subseteq A$ and $B^{p}\subseteq B$, with the following properties:
@@
A structure is countably saturated if every consistent type over a countable set is realised.
Thus ‘countably saturated’ is what is usually called ‘$\aleph_{1}$-saturated’ and should not be confused with the notion of ‘countably saturated’ as defined in e.g., [25, §2.3] (such models are countable by definition, and countably saturated models in our sense are necessarily uncountable). The reason for this change of terminology is the same as the reason why countable additivity of Lebesgue measure is rarely referred to as $\aleph_{1}$-additivity. If the language is countable, then countable saturation is equivalent to countable compactness, i.e., the requirement that every countable consistent type is realised.
If two structures are countably saturated and elementarily equivalent, then there is a $\sigma$-complete back-and-forth system between them (see [61, Theorem 16.6.4]). This implies a special case of the following theorem due to Keisler (e.g., [25, Theorem 5.1.13]).
@@
The countable saturation of metric structures of density character $\aleph_{1}$ also has fundamental effects on the automorphism groups, since it allows the construction of $2^{\aleph_{1}}$ distinct $\sigma$-complete back-and-forth systems, giving rise to $2^{\aleph_{1}}$ distinct automorphisms of the structure (see [61, Theorem 16.6.3]).
@@ 6.1.2. Borel quotients of $\mathcal{P}({\mathbb{N}})$
However, in [99] CH was used to prove that the quotients over ${\mathcal{Z}}_{0}$, ${\mathcal{Z}}_{\log}$, and even all so-called EU-ideals, are isomorphic. The intricate back-and-forth construction produced an isometric isomorphism. It took a couple of decades before this heroic achievement was put into a proper context: All quotients of $\mathcal{P}({\mathbb{N}})$ by generalised density ideals (Definition 3.5) are countably saturated metric structures by Proposition 5.11 and Theorem 6.8. This fact was used in [56, Theorem 2] to prove that under CH there are only two non-isomorphic quotients over dense density ideals, and it implies the following.
@@
Assume $\mathrm{CH}$ and let $\mathcal{I}\subseteq\mathcal{P}({\mathbb{N}})$ be a generalised density ideal. Then there are $2^{\mathfrak{c}}$ topologically nontrivial automorphisms of $\mathcal{P}({\mathbb{N}})/\mathcal{I}$.
@@ 6.1.3. Zero-dimensional remainders
In [127] Parovičenko characterised, under $\mathrm{CH}$, $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ as the only space sharing certain topological properties (such properties define a Parovičenko space, in today’s terminology). These topological properties are indeed shared by all spaces of the form $\beta X\setminus X$, where $X$ is a second countable locally compact noncompact zero-dimensional space [^18] It was proved in [172] that $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ being unique among Parovičenko spaces is equivalent to $\mathrm{CH}$.. Through Stone duality, Parovičenko’s theorem essentially states that the Boolean algebras of clopen sets of those remainders are atomless and countably saturated. Since the theory of atomless Boolean algebras is complete, all of its models are elementarily equivalent. With the definition of algebraically trivial homeomorphisms (Definition 4.8) in mind and the fact that topologically trivial automorphisms of coronas of abelian $\mathrm{C}^{*}$-algebras are dual to algebraically trivial homeomorphisms ([178, Proposition 2.7]), this implies the following.
@@
On the positive side, if $X$ is equal to a strictly increasing union of compact subsets $K_{n}$ which satisfy ($\partial K_{n}$ is the topological boundary of $K_{n}$) $\sup_{n}|\partial K_{n}|<\infty$, then $C(\beta X\setminus X)$ is countably saturated (this is a special case of [75, Theorem 2.5], and includes the case $X=\mathbb{R}$).
@@ 6.1.4. Countable saturation of the reduced products
Our next task is to review some of the applications of Theorem 6.3 in the noncommutative setting of reduced products. Ultraproducts associated with countably incomplete ultrafilters are countably saturated (see [25, Theorem 6.1.1]). This, in particular, is the case for the norm ultraproducts of $\mathrm{C}^{*}$-algebras. As for coronas of separable $\mathrm{C}^{*}$-algebras, they are not necessarily countably saturated. Notably, the Calkin algebra is not countably saturated (§6.3). However, corona algebras that are reduced products over $\operatorname{Fin}$ are countably saturated by the following result ([75], see also [61, Theorem 16.5.1]).
@@ 6.1.5. Metric Feferman–Vaught theorem
Theorem 6.8 is proved by combining the metric version of the classical Feferman–Vaught theorem with the proof of the countable saturation of ultraproducts associated with countably incomplete ultrafilters (see e.g., [61, §16.5]). In classical model theory Feferman and Vaught ([79]) gave an effective way to determine the satisfaction of formulas in the reduced products of models of the same language associated with any ideal over ${\mathbb{N}}$. Their result was generalised in [84] to metric structures. The metric Feferman–Vaught theorem essentially states that if $(A_{n})_{n\in{\mathbb{N}}}$ are $\mathcal{L}$-structures and $\mathcal{I}$ is an ideal on ${\mathbb{N}}$ then the theory of the reduced product $\prod_{n}A_{n}/\bigoplus_{\mathcal{I}}A_{n}$ can be recursively computed from the theories of $A_{n}$ and the theory of $\mathcal{P}({\mathbb{N}})/\mathcal{I}$ as a Boolean algebra.
@@
An elementary embedding from $C(K,A)$ into $A_{\infty}$ was exhibited in [63, Proposition 3.5], and the general case follows from Theorem 6.10. ∎
@@ 6.1.6. The Palyutin–Palmgren approach
The key to Palmgren’s proof is the observation that in every reduced power associated with $\operatorname{Fin}$ (or any other ideal such that the quotient Boolean algebra is atomless) every formula is equivalent to a Boolean combination of $h$-formulas. Therefore every type over such reduced product is axiomatised by a set consisting of $h$-formulas and negations of $h$-formulas. This proof was extended to prove countable saturation of reduced powers over a larger class of the so-called layered ideals in [31, Theorem 4.1].
@@ 6.1.7. Asymptotic sequence algebras
For discrete structures ‘density character’ is of course a synonym for cardinality. For more see [63, Theorem A–F].
@@ 6.2. Full saturation: The dividing line
A result analogous to (6.2.1) for ultrapowers appears in [69, Theorem 5.6], and the original proof of (6.2.1) closely followed the proof of the latter. This proof was considerably simplified by using ideas of Palyutin and Palmgren (see §6.1.6). Item (6.2.2) is proven by injecting $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ into the quotient in a gap-preserving manner. Its proof applies in a wider context, showing for example that if $M$ is a reduced product over the asymptotic density zero ideal ${\mathcal{Z}}_{0}$ (see §3.1), or any other ${\mathcal{I}}$ such that $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ is not countably saturated, whose theory is not stable then $M$ is not even countably saturated ([31, Corollary 2.3]).
@@ 6.3.1. The Fredholm index obstruction
The simplest example of the failure of countable saturation in the Calkin algebra is related to the original Brown–Douglas–Fillmore question, whether there is an automorphism of $\mathcal{Q}(H)$ that sends $\dot{s}$ (where $s$ is the unilateral shift) to its adjoint (see Question 4.16). This is really a question about orbits of unitaries in the Calkin algebra under the action of the automorphism group by conjugation. Every unitary $u$ in $\mathcal{Q}(H)$ lifts to a partial isometry (see (4.1.15)) between closed subspaces of $H$ of finite co-dimension. The difference between the co-dimensions of the domain and the range of the lift is its Fredholm index (§4.3). The function that sends $u$ to the Fredholm index of its lift is a group homomorphism from $\mathcal{U}(\mathcal{Q}(H))$ into $({\mathbb{Z}},+)$. This implies that the type of a unitary that (for example) has $2^{n}$th root for all $n$, but no $3$rd root, is consistent but not realised in $\mathcal{Q}(H)$. (The proof uses definability of the set of unitaries, in the sense of [9], see also [68, §3] to show that in $\mathcal{Q}(H)$—and in any $\mathrm{C}^{*}$-algebra—applying quantification over the unitary group to a definable predicate results in a definable predicate). By adding a variable for each $2^{n}$th root, this example provides a universal type which is consistent but not realised in $\mathcal{Q}(H)$.
@@ 6.3.2. Homotopy
By iterating the construction of Example 6.16 one obtains the following, due to N.C. Phillips ([67, Proposition 4.2]); recall the convention that $\dot{p}$ denotes a lift of $p$.
@@
There is a unital copy $A$ of the CAR algebra $\bigotimes_{\mathbb{N}}M_{2}(\mathbb{C})$ in $\mathcal{Q}(H)$ such that for every projection $p\in\mathcal{B}(H)$ with $\dot{p}\in A$, none of the tensor factors $M_{2}(\mathbb{C})$ lifts to a unital copy of $M_{2}(\mathbb{C})$ in $p\mathcal{B}(H)p$ (for more see [145]). Fix a unital copy $B$ of the CAR algebra that lifts to a unital copy of the CAR algebra in $\mathcal{B}(H)$ and an isomorphism $\Phi\colon A\to B$. The type of a unitary that implements this isomorphism is, by Example 6.16, consistent, but it is not realised in $\mathcal{Q}(H)$.
The Galois type of a tuple (i.e., its orbit under the action of the automorphism group) provides more information than its first-order type. Example 6.16 shows that in $\mathrm{C}^{*}$-algebras, the Galois type of a tuple needs to be supplemented with its homotopy type. There is an alternative way of defining the homotopy type of a tuple in a $\mathrm{C}^{*}$-algebra. For a definable set $X$ (such as the set of unitaries or the set of projections; see [68, §3]) one says that two elements of $X$ are homotopic if they can be connected by a continuous path inside $X$. This approach is used in defining K-theory groups of $\mathrm{C}^{*}$-algebras (see [137]), and we will take advantage of it in the next subsection.